What is the wavelength (in meters) of an electromagnetic waves whose frequency is 3.66 X107 Hz


1
Expert's answer
2021-01-06T05:56:00-0500

  λ = "\frac{c}{f}"

  λ = wavelength , c = speed of light = 3 × 108 , f = frequency = 3.66 × 107

  λ = "\frac{3×10^8}{3.66×10^7}"


  λ = 8.196721311 m

Or λ = 8.20 m

m